The aurora is caused when electrons and protons, moving in the earth's magnetic field of ≈5×10−5 T, collide with molecules of the atmosphere and cause them to glow. What is the radius of the circular orbit for (a) an electron with speed 1.0×106 m/s and (b) a proton with speed 7.0×104 m/s . Part A Express your answer using two significant figures.
Answer #1

R = mV/qB

=>

a) R = (9.1*10^-31*10^6)/(1.6*10^-19*5*10^-5) = 0.11 m

b) R = (1.67*10^-27*7*10^4)/(1.6*10^-19*5*10^-5) = 14.61 m